# AdvGPS: Adversarial GPS for Multi-Agent Perception Attack

### 30 Jan 2024 v1

@@ Abstract
The multi-agent perception system collects visual data from sensors located on various agents and leverages their relative poses determined by GPS signals to effectively fuse information, mitigating the limitations of single-agent sensing, such as occlusion. However, the precision of GPS signals can be influenced by a range of factors, including wireless transmission and obstructions like buildings. Given the pivotal role of GPS signals in perception fusion and the potential for various interference, it becomes imperative to investigate whether specific GPS signals can easily mislead the multi-agent perception system. To address this concern, we frame the task as an adversarial attack challenge and introduce AdvGPS, a method capable of generating adversarial GPS signals which are also stealthy for individual agents within the system, significantly reducing object detection accuracy. To enhance the success rates of these attacks in a black-box scenario, we introduce three types of statistically sensitive natural discrepancies: appearance-based discrepancy, distribution-based discrepancy, and task-aware discrepancy. Our extensive experiments on the OPV2V dataset demonstrate that these attacks substantially undermine the performance of state-of-the-art methods, showcasing remarkable transferability across different point cloud based 3D detection systems. This alarming revelation underscores the pressing need to address security implications within multi-agent perception systems, thereby underscoring a critical area of research.
@@ II Related Work
Deployment of Cooperative Perception. Multi-agent perception systems offer numerous advantages but are accompanied by challenges such as localization errors, communication latency, adversarial attacks, and lossy communication. These development challenge can easily diminish the benefits of collaborations [9, 6]. Several strategies have been proposed to enhance robustness. For instance, V2X-ViT employs a Vision Transformer to mitigate GPS localization errors and sensing information delays during intermediate collaboration [6]. To address localization errors, [10] proposes a pose regression module that learns correction parameters to predict accurate relative transformations from noisy data. The Lossy Communication-aware Repair Network (LCRN) [4] is introduced to address packet loss issues in communication. A model-agnostic framework [11] has also been proposed to handle model heterogeneity in collaborative perception. A Multi-agent Perception Domain Adaption (MPDA) framework [12] is proposed for cooperative perception to bridge the domain gap for multi-agent perception. Nevertheless, V2V cooperative perception performance relies heavily on GPS signals, which might be susceptible to real-world attacks. This challenge might lead to a significant degradation in perception performance. This paper investigates GPS attacks in the context of 3D object detection tasks within V2V cooperative perception.
@@ IV-A Overview
where $\mathbf{F}_{cav\rightarrow ego}=\text{Encoder}(\mathbf{T}_{cav\rightarrow ego}\mathbf{P}_{cav})$ and $\mathbf{F}_{ego}=\text{Encoder}(\mathbf{P}_{ego})$ denote the features of the two point clouds $\mathbf{P}_{cav\rightarrow ego}$ and $\mathbf{P}_{ego}$, $\text{Fusion}(\cdot)$ is a network to fuse the two features, and $\varphi(\cdot)$ is the header for object detection. It aims to exploit adversarial attacks on GPS signals to deceive any point cloud based 3D detection models in a black-box setting. To achieve this, we calculate an imperceptible noise-like perturbation under the guidance of a classic 3D object detection model and add it to the original CAV GPS pose to obtain an AdvGPS pose. This corrupted AdvGPS pose can then mislead any other detection models in a black-box setting. Our AdvGPS attack perturbation is stealthy to be confused as normal GPS signal errors. To enhance the success rates of attacks in black-box scenarios, we consider three objectives to optimize the adversarial GPS: appearance-based discrepancy, distribution-based discrepancy, and detection task-aware discrepancy. We define the optimization over the adversarial GPS signal $\widehat{G}_{cav}$ as follows:
@@ IV-E Implementation Details
To implement our AdvGPS for V2V cooperative perception attack, we follow the general adversarial attack procedure: ❶ We set the additive perturbation $\mathbf{g}$ to $\widehat{G}_{cav}$ and obtain the original CAV pose to calculate the HTM. ❷ Based on the HTM, we calculate $D_{app}$ for the projected point cloud. ❸ We feed the projected point cloud into an encoder to generate the intermediate feature and calculate $D_{dist}$. ❹ The intermediate features are then inputted into the fusion network to calculate $D_{task}$. ❺ We perform back propagation and compute the gradients of $\mathbf{g}$ with respect to the loss functions. ❻ The sign of the gradients is used to update additive perturbation $\mathbf{g}$ by multiplying them with a step size, then adding $\mathbf{g}$ to the $\widehat{G}_{cav}$. ❼ We calculate a new synthesized $\mathbf{T}_{cav\rightarrow ego}$ via Eq. 2 and repeat steps ❷ to ❻ for a defined number of iterations. We set the number of iterations to 10. To conceal within the range of actual GPS errors, we follow the statistics presented in [21], which indicates that the real-world GPS signal’s average localization error, height error, and heading error are $1.118$ meters, $1.395$ meters and $0.141$ degrees, respectively, so we set $\epsilon_{x,y}$, $\epsilon_{z}$, and $\epsilon_{\theta_{x},\theta_{y},\theta_{z}}$ as 1.118, 1.395, and 0.141, respectively.
@@ V Experiments
Figure 4: 3D detection visualization on attacking V2V model CoBEVT [8]. Blue and red 3D bounding boxes represent ground truth and prediction. Purple arrows: detection errors. Green point cloud: by ego vehicle. Other-color point clouds: projected to ego coordinate by nearby attacked CAVs. The small shift of zoomed area verifies that AdvGPS attack is stealthy.
@@ V-A Experimental Setups
Evaluation metrics: We evaluate the performance of our proposed framework by the final 3D vehicle detection accuracy. Following [1, 6], we set the evaluation range as $x\in[-140,140]$ meters, $y\in[-40,40]$ meters, where all CAVs are included in this spatial range, whose number is in the range of $[1,5]$ in the experiment. We measure the accuracy with Average Precision (AP) at the Intersection-over-Union (IoU) threshold of 0.5.
Attack settings: To evaluate the effectiveness of our AdvGPS attack on the V2V perception pipeline, we compare it with three commonly used attack methods on CAVs’ GPS poses: Projected Gradient Descent (PGD) [24], Fast Gradient Sign Method (FGSM) [22], and Iterative Fast Gradient Sign Method (IFGSM) [23]. Additionally, we include two current GPS spoofing methods: Random Bias Attack (RBA) [5] and Position Altering Attack (PAA) [18]. FGSM, IFGSM, PGD, and PAA are gradient-based attack methods, following the implementation settings described in Section IV-E. RBA can be implemented by applying uniformly distributed random bias within the same range settings ($\epsilon_{x,y}$, $\epsilon_{z}$ and $\epsilon_{\theta_{x},\theta_{y},\theta_{z}}$) as in Section IV-E. We use the Adam optimizer [27], and all models are executed on two RTX 3090 GPUs. We conduct both black-box and white-box attacks to assess the vulnerability and transferability of the cooperative perception methods. These attack scenarios are described as follows:
@@
  • Black-Box Attack (B.b.): We employ VoxelNet [20] as the encoder and a simple self-attention model as the feature fusion network. We save the adversarial GPS information of all attacked CAVs across the entire frame duration to be used for attacking PointPillar [26] encoder and more advanced fusion networks based cooperative perception methods.
@@ V-B Quantitative Evaluation
Sensitivity analysis of GPS parameters: We have conducted a sensitivity analysis for each parameter of CAV poses $\widehat{G}_{cav}$ in Table II. Even when operating within the boundaries of typical GPS errors, both the PAA [18] and our AdvGPS methods remain effective in significantly reducing object detection accuracy for two ViT-based models, namely V2X-ViT [6] and CoBEVT [8]. Intuitively, among all six parameters, attacking the positional parameters $x$ and $y$ of CAVs’ poses is the most efficient approach within a small range of GPS pose changes, highlighting their sensitivity within the V2V cooperative perception system. Furthermore, when we set the positional parameters $[x,y,z]$ to have maximum GPS bias $[1.118,1.118,1.395]$ meters of the actual measured GPS errors in real world [21] and attack Attfuse [1], we obtained a AP $65.1\%$ at IoU 0.5, close to the performance by RBA [5], which has less degradation than our AdvGPS.
Ablation study: As illustrated in Table III, our comprehensive ablation study reveals that all three components within our AdvGPS attack framework contribute significantly to the degradation of detection performance on CoBVET [8]. In the context of the Black-Box Attack setting, the inclusion of $D_{app}$, $D_{dist}$, and $D_{task}$ results in performance degradation by $48.1\%$, $37.2\%$, and $27.0\%$, respectively. This nuanced analysis emphasizes the effectiveness of our design. It is evident that our proposed AdvGPS attacks significantly undermine the performance of state-of-the-art V2V cooperative perception methods.

### 20 Feb 2024 v2

@@ Abstract
The multi-agent perception system collects visual data from sensors located on various agents and leverages their relative poses determined by GPS signals to effectively fuse information, mitigating the limitations of single-agent sensing, such as occlusion. However, the precision of GPS signals can be influenced by a range of factors, including wireless transmission and obstructions like buildings. Given the pivotal role of GPS signals in perception fusion and the potential for various interference, it becomes imperative to investigate whether specific GPS signals can easily mislead the multi-agent perception system. To address this concern, we frame the task as an adversarial attack challenge and introduce AdvGPS, a method capable of generating adversarial GPS signals which are also stealthy for individual agents within the system, significantly reducing object detection accuracy. To enhance the success rates of these attacks in a black-box scenario, we introduce three types of statistically sensitive natural discrepancies: appearance-based discrepancy, distribution-based discrepancy, and task-aware discrepancy. Our extensive experiments on the OPV2V dataset demonstrate that these attacks substantially undermine the performance of state-of-the-art methods, showcasing remarkable transferability across different point cloud based 3D detection systems. This alarming revelation underscores the pressing need to address security implications within multi-agent perception systems, thereby underscoring a critical area of research. The code is available at https://github.com/jinlong17/AdvGPS.
@@ II Related Work
Deployment of Cooperative Perception. Multi-agent perception systems offer numerous advantages but are accompanied by challenges such as localization errors, communication latency, adversarial attacks, and lossy communication. These development challenge can easily diminish the benefits of collaborations [9, 6]. Several strategies have been proposed to enhance robustness. For instance, V2X-ViT employs a Vision Transformer to mitigate GPS localization errors and sensing information delays during intermediate collaboration [6]. To address localization errors, [10] proposes a pose regression module that learns correction parameters to predict accurate relative transformations from noisy data. The Lossy Communication-aware Repair Network (LCRN) [4] is introduced to address packet loss issues in communication. A Multi-agent Perception Domain Adaption (MPDA) framework [11] is proposed for cooperative perception to bridge the domain gap for multi-agent perception. Nevertheless, V2V cooperative perception performance relies heavily on GPS signals, which might be susceptible to real-world attacks. This challenge might lead to a significant degradation in perception performance. This paper investigates GPS attacks in the context of 3D object detection tasks within V2V cooperative perception.
@@ IV-A Overview
where $\mathbf{F}_{cav\rightarrow ego}=\text{Encoder}(\mathbf{T}_{cav\rightarrow ego}\mathbf{P}_{cav})$ and $\mathbf{F}_{ego}=\text{Encoder}(\mathbf{P}_{ego})$ denote the features of the two point clouds $\mathbf{P}_{cav\rightarrow ego}$ and $\mathbf{P}_{ego}$, $\text{Fusion}(\cdot)$ is a network to fuse the two features, and $\varphi(\cdot)$ is the regression and detection header for object detection [1]. It aims to exploit adversarial attacks on GPS signals to deceive any point cloud based 3D detection models in a black-box setting. We calculate an imperceptible noise-like perturbation under the guidance of a classic 3D object detection model and add it to the original CAV GPS pose to obtain an AdvGPS pose. This corrupted AdvGPS pose can then mislead any other detection models in a black-box setting. Our AdvGPS attack perturbation is stealthy to be confused as normal GPS signal errors. To enhance the success rates of attacks in black-box scenarios, we consider three objectives to optimize adversarial GPS: appearance-based discrepancy, distribution-based discrepancy, and detection task-aware discrepancy. We define the optimization over the adversarial GPS signal $\widehat{G}_{cav}$ as:
@@ IV-E Implementation Details
To implement our AdvGPS for V2V cooperative perception attack, we follow the general adversarial attack procedure: ❶ We set the additive perturbation $\mathbf{g}$ to $\widehat{G}_{cav}$ and obtain the original CAV pose to calculate the HTM. ❷ Based on the HTM, we calculate $D_{app}$ for the projected point cloud. ❸ We feed the projected point cloud into an encoder to generate the intermediate feature and calculate $D_{dist}$. ❹ The intermediate features are then inputted into the fusion network to calculate $D_{task}$. ❺ We perform back propagation and compute the gradients of $\mathbf{g}$ with respect to the loss functions. ❻ The sign of the gradients is used to update additive perturbation $\mathbf{g}$ by multiplying them with a step size, then adding $\mathbf{g}$ to the $\widehat{G}_{cav}$. ❼ We calculate a new synthesized $\mathbf{T}_{cav\rightarrow ego}$ via Eq. 2 and repeat steps ❷ to ❻ for a defined number of iterations. We set the number of iterations to 10. To conceal within the range of actual GPS errors, we follow the statistics presented in [20], which indicates that the real-world GPS signal’s average localization error, height error, and heading error are $1.118$ meters, $1.395$ meters and $0.141$ degrees, so we set $\epsilon_{x,y}$, $\epsilon_{z}$, and $\epsilon_{\theta_{x},\theta_{y},\theta_{z}}$ as 1.118, 1.395, and 0.141 respectively to avoid significant offsets.
@@ V Experiments
Figure 4: 3D detection visualization on attacking V2V model CoBEVT [8]. Blue and red 3D bounding boxes represent ground truth and prediction. Purple arrows: detection errors. Green point cloud: by ego vehicle. Other-color point clouds: projected to ego coordinate by nearby attacked CAVs.
@@ V-A Experimental Setups
Evaluation metrics: We assess the final 3D vehicle detection accuracy for our proposed framework performance. Similar to prior works [1, 6], we set $x\in[-140,140]$ meters, $y\in[-40,40]$ meters as the evaluation range, where all CAVs (in the range of $[1,5]$) are included in this spatial range. We employ Average Precision (AP) at the Intersection-over-Union (IoU) threshold of 0.5 as the accuracy metric.
Attack settings: To evaluate the effectiveness of our AdvGPS attack on the V2V perception pipeline, we compare it with three commonly used attack methods on CAVs’ GPS poses: Projected Gradient Descent (PGD) [23], Fast Gradient Sign Method (FGSM) [21], and Iterative Fast Gradient Sign Method (IFGSM) [22]. Additionally, we include two current GPS spoofing methods: Random Bias Attack (RBA) [5] and Position Altering Attack (PAA) [17]. FGSM, IFGSM, PGD, and PAA are gradient-based attack methods, following the same implementation settings described in Section IV-E. RBA can be implemented by applying uniformly distributed random bias within the same range settings ($\epsilon_{x,y}$, $\epsilon_{z}$ and $\epsilon_{\theta_{x},\theta_{y},\theta_{z}}$) as in Section IV-E. We use the Adam optimizer [26], and all models are executed on two RTX 3090 GPUs. We conduct both black-box and white-box attacks to assess the vulnerability and transferability of the cooperative perception methods. These attack scenarios are described as:
@@
  • Black-Box Attack (B.b.): We employ VoxelNet [19] as the encoder and a naive self-attention model as the feature fusion network in training. We save the adversarial GPS information of all attacked CAVs across the entire frame duration to be used for attacking the unseen PointPillar [25] encoder and other state-of-the-art fusion methods of cooperative perception.
@@ V-B Quantitative Evaluation
Sensitivity analysis of GPS parameters: We have conducted a sensitivity analysis for each parameter of CAV poses $\widehat{G}_{cav}$ in Table II. Even when operating within the boundaries of typical GPS errors, both the PAA [17] and our AdvGPS methods remain effective in significantly reducing object detection accuracy for two ViT-based models, namely V2X-ViT [6] and CoBEVT [8]. Intuitively, among all six parameters, attacking the positional parameters $x$ and $y$ of CAVs’ poses is the most efficient approach within a small range of GPS pose changes, highlighting their sensitivity within the V2V cooperative perception system. Furthermore, when we set the positional parameters $[x,y,z]$ to have maximum GPS bias $[1.118,1.118,1.395]$ meters of the actual measured GPS errors in real world [20] and attack Attfuse [1], we obtained a AP $65.1\%$ at IoU 0.5, close to the performance by RBA [5], which has less degradation than our AdvGPS. It demonstrates that the gradient optimization based attack methods are better than the non-gradient attack.
Ablation study: As shown in Table III, our comprehensive ablation study reveals that all three components within our AdvGPS attack framework contribute significantly to the degradation of detection performance on CoBVET [8]. In the context of the Black-Box Attack setting, the inclusion of $D_{app}$, $D_{dist}$, and $D_{task}$ results in performance degradation by $48.1\%$, $37.2\%$, and $27.0\%$, respectively. This nuanced analysis emphasizes the effectiveness of our design. It is evident that our proposed AdvGPS attacks significantly undermine the performance of state-of-the-art V2V cooperative perception methods.